1.     A ball is thrown horizontally from a high window at 6 m/s and reaches the ground after 2 s.  Calculate:

a)         the horizontal distance travelled

b)        the vertical speed at impact.


2.     A stone thrown horizontally from a cliff lands 24 m out from the cliff after 3 s. Find:

a.     the horizontal speed of the stone 

b.     the vertical speed at impact.


3.     A ball is projected horizontally at 15 m/s from the top of a vertical cliff. It reaches the ground 5 s later.

a.     Calculate the horizontal distance it travels in 5seconds

b.     What was its vertical speed just before it hit the ground?





1
Expert's answer
2021-04-20T16:49:38-0400

"\text {body formulas thrown horizontally:}"

"V_x = const ; S_x=V_x*t"

"V_y =g*t;S_y=\frac{gt^2}{2}"


"1. V_x=6m\/s;t=2s"

"a)S_x = V_x*t = 6*2 =12m"

"b)V_y =g*t = 9.8*2 = 19.6 m\/s"


"2.S_x =24m;t=3s"

"a)V_x=\frac{S_x}{t}=\frac{24}{3}=8m/s"

"b)V_y =g*t = 9.8*3 = 29.4 m\/s"


"3.V_x=15 m\/s;t=5s"

"a)S_x = V_x*t = 15*5 =75m"

"b)V_y =g*t = 9.8*5 = 48m\/s"


Answer: 1. a)12 m b)19.6 m/s

2. a)8 m/s b)29.4 m/s

3. a)75 m b)48m/s